OpenStudy (anonymous):

In one week a gas station sells 2,570 gallons of gas. On average, how many quarts did they sell each day during that week?

OpenStudy (anonymous):

Ok so there are 4 quarts in a gallon and there are 7 days a week. \[2570gallons*\frac{ 1week }{ 7days }*\frac{ 1gallon }{ 4quarts }\]So when figuring out just do\[2570*\frac{ 1 }{ 7 }*\frac{ 1 }{ 4 }=\frac{ 2570 }{ 28 }quarts per day\]

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

Yep